What is the Molarity of Fe(CN)6^4- in a saturated solution of Ag4[Fe(CN)6]? (Ksp = 1.6×10^-14)??

Grade:12

1 Answers

Aaditee
5 years ago
According to the theory:-
{Ag}{Fe(CN)6}=Ksp
x*4^4*x^4=Ksp
256x^5=1.6*10^-41
x^5=6.25*10^-44
Therefore x=2.3*10^-9